OpenStudy (anonymous):

4/(m+2) - (3m)/(m-3)

OpenStudy (anonymous):

=4(m-3)/(m+2)(m-3) - 3m(m+2)/(m+2)(m-3) =4m-12-(3m^2+6m)/(m+2)(m-3) =-3m^2-2m-12/(m+2)(m-3) hope this helps!